# Acceleration deforms exponential decays into power laws

### v4

@@ Abstract
An exponentially decaying accelerated system looks as if its decay was of a generalized power-law type, provided relativistic time dilation in the detector and retardation of the emitted signal are taken into account. The same mathematical formula is found in generalizations of the Zipf-Mandelbrot law in quantitative linguistics, and in dynamics of re-association of folded proteins.
@@ I Watching a decay
Assume an unstable system propagating along a world-line $r\mapsto x^{\mu}(r)$ in Minkowski space decays exponentially in its own reference frame. The survival probability at proper time $r$ is given by $p(r)=e^{-r/r_{0}}$. Here, $r\geq 0$ is the proper time computed along the world-line and $r_{0}$ is the mean lifetime (we measure proper time in units of length). Now, consider another world-line, $s\mapsto y^{\mu}(s)$, describing a detector which absorbs at $y^{\mu}(s)$ the light signal emitted at $x^{\mu}(r)$. Here, $s\geq 0$ is the proper time of the detector, computed along the detector’s world-line. The difference $y^{\mu}(s)-x^{\mu}(r)$ is a future-pointing null vector. We also assume that at the moment the decay begins, at $r=s=0$, both objects are at the same point in space-time, $x^{\mu}(0)=y^{\mu}(0)$. If the decay takes place at proper time $r$, it will be detected at proper time $s=s(r)$ which depends on both world-lines (Fig. 1).
@@
The problem is nontrivial in that accelerated sources or detectors may lead to event horizons. In particular, a particle that decays exponentially in its rest frame may approach the event horizon of the detector in a finite time $r_{\rm max}$, but the detector will have to wait forever to see the source cross the horizon if $s(r_{\rm max})=\infty$. Accordingly, $P(s)$ cannot be exponential because $P(\infty)\neq 0$.
The probability of non-detection, $P(s)$, equals the probability of non-emission, when signal retardation and time dilation in the detector are taken into account. Hence, we assume
@@
The exponential decay is then replaced by an exact power law, but the event horizon disappears.
The replacement of an exponential law by a power law is a purely relativistic effect. It should not be confused with other possible reasons for deviations from a simple exponential formula one finds in experiment. We know, for example, that the spontaneous emission of an excited atomic state is exponential only approximately, but this is not what we are concerned with here. The formula we derive below will have its analogue for any form of the decay, and the result will be again given by (1). Systems accelerating in more complicated ways will lead to more complicated forms of $s(r)$, still satisfying our differential equation for $ds(r)/dr$. Our derivation is generally valid for any motion of sources and detectors.
Somewhat unexpectedly, the probability given by (4) (especially when written as (75)) turns out to have exactly the form postulated over two decades ago by Tsallis, Bemski and Mendes [1] in their analysis of re-association of folded proteins [2, 3]. Moreover, inspired by [1], Montemurro [4] showed that the same probability function can be used to adequately model observable deviations from the Zipf-Mandelbrot law in quantitative linguistics. All these links with the problem discussed in the present paper were completely unexpected and hard to anticipate. It should be stressed, though, that neither [1] nor [4] were capable of deriving (4) and (75) from first principles — rather, it was an educated guess based on the structure of the data and an ad hoc modification of some differential equations.
@@
We begin in Section II with determining the form of the map $r\mapsto s(r)$ which relates the proper time of emission with the one of the detection. We derive a differential equation satisfied by $s(r)$ once the explicit forms of $x^{\mu}(r)$ and $y^{\mu}(s)$ are given. In Section III we restrict our analysis to the concrete case of sources and detectors that move with constant accelerations. The two world-lines cross twice: at the initial condition and then once again, ultimately separating into their asymptotic forms at infinity. The two stages, between the crossing points and behind the second of them, are qualitatively different and have to be treated independently. Once we have obtained the explicit form of $s(r)$ we are able to write our final formula for $P(s)$, the task completed in Section IV. Finally, in Section V we briefly discuss implications of our analysis for the distinction between relativistic clocks and watches.
@@ II Relation between proper times
Assume the three world-lines are restricted to some $t$-$z$ plane of the Minkowski space,
@@
We begin with establishing a relation between $\sigma^{2}=\Delta y_{\mu}(s)\Delta y^{\mu}(s)$, $\rho^{2}=\Delta x_{\mu}(r)\Delta x^{\mu}(r)$, and $\Delta q$. All the null vectors in Fig. 1 and Fig. 3 are parallel. In particular,
@@
The situation changes behind the second crossing point (Fig. 5), where we find that $\gamma(s)$ is negative, $\gamma(s)=-|\gamma(s)|$ but the sign in $\sigma_{\pm}$ is positive. Hence,
@@ IV Effective non-exponential decay
in order to fit the protein folding data from [2, 3], and employed in [4] in fitting Zipf-law data from 36 plays by Shakespeare and 56 books by Dickens. Although there is no reason to believe that there are any links between the statistics of muon decays and frequencies of words used by Shakespeare, it is nevertheless intriguing that the same non-evident functional dependence on parameters is found.
@@
V Relativistic clocks and watches
Our result, as based on special relativity, cannot be in conflict with the standard tests of time dilation [12, 13]. The experiments were always designed to measure directly the decay, and not the decay products. Our conclusions may thus have some consequences for astrophysical or accelerator observations of atomic or particle lifetimes, but at this stage we cannot say anything more concrete.
Yet, it should be mentioned that Fig. 6 suggests another perspective on the subject of the paper. Namely, let us note that both $r$ and $s(r)$ are proper times measured by clocks propagating along the two world-lines. $s(r)$ is the proper time registered by the detector at the moment it detects (i.e. observes) the light signal emitted at proper time $r$ (as measured by the clock of the source). The clock of the source thus plays the role of a watch. An observer propagating along $s\mapsto y^{\mu}(s)$ sees at his proper time $s(r)$ the time $r$ as it appears on the clock of the source located at $x^{\mu}(r)$. The relation between the time $r$ of emission and the time $s(r)$ of observation is one-to-one. Therefore if we watch a decaying accelerating system, what we see is not the exponential decay but a generalized power law.

### v5

# Acceleration deforms exponential decays into generalized Zipf-Mandelbrot laws

@@ Abstract
An exponentially decaying system looks as if its decay was a generalized power or double-exponential law, provided one takes into account the relativistic time dilation in a detector, the delay of the emitted signal, and the accelerations of both the source and the detector. The same mathematical formula can be found in generalizations of the Zipf-Mandelbrot law in quantitative linguistics and in the dynamics of ligand binding in heme proteins. The effect is purely kinematic and is not related to the various dynamic phenomena that can accompany accelerated motion of sources or detectors. The procedure used can also be seen as a form of clock synchronization near an event horizon.
@@ I Watching a decay
Assume an unstable system propagating along a world-line $r\mapsto x^{\mu}(r)$ in Minkowski space decays exponentially in its comoving reference frame [1, 2]. The survival probability at proper time $r$ is given by $p(r)=e^{-r/r_{0}}$. Here, $r\geq 0$ is the proper time computed along the world-line and $r_{0}$ is the mean lifetime (we measure proper time in units of length). Now, consider another world-line, $s\mapsto y^{\mu}(s)$, describing a detector which absorbs at $y^{\mu}(s)$ the light signal emitted at $x^{\mu}(r)$. Here, $s\geq 0$ is the proper time of the detector, computed along the detector’s world-line. The difference $y^{\mu}(s)-x^{\mu}(r)$ is a future-pointing null vector. We also assume that at the moment the decay begins, at $r=s=0$, both objects are at the same point in space-time, $x^{\mu}(0)=y^{\mu}(0)$. If the decay takes place at proper time $r$, it will be detected at proper time $s=s(r)$ which depends on both world-lines (Fig. 1).
@@
The problem is nontrivial in that accelerated sources or detectors may lead to event horizons. In particular, a particle that decays exponentially in its rest frame may approach the event horizon of the detector in a finite time $r_{\rm max}$, but the detector will have to wait forever to see the source cross the horizon if $s(r_{\rm max})=\infty$. Accordingly, the survival probability of the detector, $P(s)$, cannot be exponential because $P(\infty)\neq 0$.
Assuming an ideal detector, the probability of non-detection equals the probability of non-emission, provided signal retardation and time dilation in the detector are taken into account:
@@
The exponential decay is then replaced by an exact power law (of a Zipf-Mandelbrot type), but the event horizon disappears.
At the other extreme is the case of an accelerating detector and a non-accelerating source,
$\lim_{a\to 0}P(s)$ $=$ $\left\{\begin{array}[]{rl}\exp\left(-\frac{1}{|b|r_{0}}(U_{0}-|\bm{U}|)(V_{0}-|\bm{V}|)(e^{|b|s/c^{2}}-1)\right),&0\leq s\leq s_{1},\\ \exp\left(-\frac{1}{|b|r_{0}}(U_{0}+|\bm{U}|)(V_{0}+|\bm{V}|)(1-e^{-|b|s/c^{2}})\right),&s\geq s_{1},\end{array}\right.$ (11)
which is a Gumbel-type double-exponential distribution.
Formulas (8) and (11) show that the two limits commute, leading to a Doppler-type formula,
$\lim_{a\to 0}\lim_{b\to 0}P(s)$ $=$ $\lim_{b\to 0}\lim_{a\to 0}P(s)$ (12)
$=$ $\exp\left(-\frac{s}{r_{0}}\frac{U_{0}-|\bm{U}|}{V_{0}+|\bm{V}|}\right)=\exp\left(-\frac{s}{r_{0}}\frac{V_{0}-|\bm{V}|}{U_{0}+|\bm{U}|}\right),$ (13)
for any $s\geq 0$ (since $s_{1}\to\infty$).
The discussed deformation of the exponential law is a purely kinematic effect caused by the relativistic time dilation in the detector and the finite speed of signal propagation. It should not be confused with other possible reasons for deviations from a simple exponential formula found in experiments.
For example, we know that spontaneous decays of unstable quantum states are only approximately exponential [3, 4, 5, 6, 7] (see [8] for a recent review). Moreover, even classical accelerated systems are subject to additional forces and potentials that can affect their lifetime, as any pyrotechnist knows well. In quantum field theory models of particle decay, acceleration leads to additional phenomena of the Fulling-Davies-Unruh type [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. However, all these effects refer to the level of $p(r)$, while what we are discussing is a result of a non-trivial $s(r)$. The latter is not affected by the form $p(r)$, but by the relation between $x^{\mu}(r)$ and $y^{\mu}(s)$.
The formula we derive below will have its analogue for any form of the decay, and the result will be again given by (1). Our derivation is generally valid for basically any motion of sources and detectors. Systems accelerating in more complicated ways will lead to more complicated forms of $s(r)$, still satisfying our differential equation for $ds(r)/dr$ (see, however, the remarks in Sec. II and Sec. V on the role of the $t$-$z$ hyperplane).
Somewhat unexpectedly, the probability given by (4) (especially when written as (84)) turns out to have exactly the form postulated over two decades ago by Tsallis, Bemski and Mendes [19] in their analysis of ligand binding dynamics [20, 21]. Moreover, inspired by [19], Montemurro [22] showed that the same probability function can be used to adequately model observable deviations from the Zipf-Mandelbrot law in quantitative linguistics. All these links with the problem discussed in the present paper were completely unexpected and hard to anticipate. It should be stressed, though, that neither [19] nor [22] were capable of deriving (4) and (84) from first principles — rather, it was an educated guess based on the structure of the data and an ad hoc modification of some differential equations.
@@
We begin in Section II with determining the form of the map $r\mapsto s(r)$ which relates the proper time of emission with the one of the detection. We derive a differential equation satisfied by $s(r)$ once the explicit forms of $x^{\mu}(r)$ and $y^{\mu}(s)$ are given. In Section III we restrict our analysis to the concrete case of sources and detectors that move with constant accelerations. The two world-lines cross twice: at the initial condition and then once again, ultimately separating into their asymptotic forms at infinity. The two stages, between the crossing points and behind the second of them, are qualitatively different and have to be treated independently. Once we have obtained the explicit form of $s(r)$ we are able to write our final formula for $P(s)$, the task completed in Section IV. Finally, in Section V we briefly discuss implications of our analysis for the problem of clock synchronization in neighborhoods of event horizons.
@@ II Relation between proper times
Assume the three world-lines shown in Fig. 1 are restricted to some $t$-$z$ plane of the Minkowski space,
@@
(14) and (15) describe, respectively, the source and the detector, whereas (16) is an auxiliary world-line introduced just for convenience and describes the origin $\bm{z}(q)=(0,0,0)$ of the rest frame.
@@
We begin with establishing a relation between $\sigma^{2}=\Delta y_{\mu}(s)\Delta y^{\mu}(s)$, $\rho^{2}=\Delta x_{\mu}(r)\Delta x^{\mu}(r)$, and $\Delta q$. The null (i.e. light-like) vectors in Fig. 3 are parallel, in consequence of restricting the motions to the $t$-$z$ hyperplane in the Minkowski space (light-like world-lines in $(1+1)$-dimensional Minkowski $t$-$z$ space are given by $ct=\pm z+\textrm{const}$, and all such lines are either parallel or perpendicular). In particular,
@@
Denoting $r=s_{a}$, $s=s_{b}$, $\gamma(s)=\gamma_{a}(s_{a})$, $\chi(r)=\gamma_{b}(s_{b})$, we can rewrite (34) as
$\frac{ds_{a}}{ds_{b}}$ $=$ $e^{\gamma_{a}(s_{a})-\gamma_{b}(s_{b})}=\beta{{}_{a}}{{}_{b}},$ (35)
exhibiting the cocycle property,
$\beta{{}_{a}}{{}_{b}}\beta{{}_{b}}{{}_{c}}=\beta{{}_{a}}{{}_{c}},$ (36)
of the world-lines and their proper-time parametrizations, i.e. the auxiliary rest-frame world-line in Fig. 1 can be replaced by any world-line passing through the two crossing points.
An analogous analysis applies to the section behind the second crossing point (Fig. 5), where we find that $\gamma(s)$ is negative, $\gamma(s)=-|\gamma(s)|$ but the sign in $\sigma_{\pm}$ is positive. Hence,
@@
(39) possess the same cocycle property (36) as the world-lines connecting the two crossing points.
@@ IV Effective non-exponential decay
in order to fit the molecular data from [20, 21], and employed in [22] in fitting the Zipf-law data from 36 plays by Shakespeare and 56 books by Dickens. Although there is no reason to believe that there are any links between the statistics of muon decays and frequencies of words used by Shakespeare, it is nevertheless intriguing that the same non-evident functional dependence on parameters is found.
@@
V Final remarks
The effect discussed is purely kinematic. We have focused on the exponential case and uniform accelerations because of the simplicity of the resulting formulas, but our results apply in principle to any decay and any motion of the sources and detectors. For example, considering various quantum and field theoretical phenomena that occur as a result of accelerations [13, 14, 15, 16, 17, 18], one expects modifications of $p(r)=e^{-r/r_{0}}$, but in any case the final result will be given by $P(s)=p(r(s))$, provided of course that all signals emitted at $x^{\mu}(r)$ are detected at $y^{\mu}(s)$.
Among the assumptions that simplify the argument is that $x^{\mu}(r)$ and $y^{\mu}(s)$ are confined to the same $t$-$z$ hyperplane for all $r$ and $s$. Source and detector thus move along the same straight line in space, making the problem effectively one-dimensional. If one relaxes this condition and considers more general motions, it would no longer be true that $y^{\mu}(s)-x^{\mu}(r)$ and $y^{\mu}(s+\Delta s)-x^{\mu}(r+\Delta r)$ are parallel for any $\Delta s$ and $\Delta r$. The simplicity of the argument would be lost. In this sense, our formulas are not the most general.
Finally, it should be mentioned that Fig. 6 suggests yet another perspective on the subject of the paper. Namely, let us note that both $r$ and $s(r)$ are proper times measured by clocks propagating along the two world-lines. $s(r)$ is the proper time registered by the detector at the moment it detects (i.e. observes) the light signal emitted at proper time $r$ (as measured by the clock of the source). An observer propagating along $s\mapsto y^{\mu}(s)$ sees at his proper time $s(r)$ the time $r$ as it appears on the clock of the source located at $x^{\mu}(r)$. The relation between the time $r$ of emission and the time $s(r)$ of observation is one-to-one. The equivalence $r\leftrightarrow s$ may be regarded as a procedure of clock synchronization in a neighborhood of an event horizon, a problem of great importance for quantum cryptography, quantum teleportation, and — more generally — the studies of entanglement in non-trivial space-times [30].